OpenStudy (jenniferjuice):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (ivettef365):

you can use the formula y = mx + b where m=slope and b= y-int

OpenStudy (anonymous):

you need to change the equation to slope intercept form

OpenStudy (ivettef365):

isolate the y on equation 5x-3y=-6

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (jenniferjuice):

slope = 5 / 3; y-intercept = (0, 2) slope = -5; y-intercept = (0, -6) slope = - 3 / 5; y-intercept = (0, 2) slope = 5 / 3; y-intercept = (0, -6)

OpenStudy (texaschic101):

you having trouble ?

OpenStudy (texaschic101):

5x - 3y = -6 first subtract 5x from each side

OpenStudy (anonymous):

If you change the equation from standard form to slope intercept you will get your answer. do you know how?

OpenStudy (jenniferjuice):

-3y=-6-5x ?

OpenStudy (anonymous):

continue trying to isolate the y.

OpenStudy (texaschic101):

-3y = - 5x - 6 now divide the equation by -3 to get y by itself

OpenStudy (anonymous):

undo the multiplication

OpenStudy (anonymous):

when dividing seperate each term. like so